Application of Quantifier Elimination Theory to Robust Multi-objective

Feedback Design

P. Dorato, W. Yang, and C. Abdallah Department of Electrical and Computer Engineering

University of New Mexico Albuquerque, NM 87131-1356 peter@eece.unm.edu

Abstract

It is shown that some important feedback design problems for systems with multiple design objectives and large parameter variations can be reduced to systems of polynomial inequalities with logic quantifiers. Quantifier elimination (QE) theory can then be used to obtain quantifier-free formulas which define the range of acceptable design parameters. A feedback design example is considered, with the constraints that the closed-loop system be robustly stable and that several design specifications in the frequency domain be met. The QE software package QEPCAD is explored for this particular example.